OpenStudy (anonymous):

Can someone please solve 3^(x-5)=9

OpenStudy (lgbasallote):

9 = 3^2 3^(x-5) = 3^2 x - 5 = 2

OpenStudy (lgbasallote):

got that?

OpenStudy (lgbasallote):

do you get the logic?

Parth (parthkohli):

As I said the last time, get the same base. \( \color{Black}{\Rightarrow 3^{x - 5} = 3^2 }\) Now you may equate the exponents. \( \color{Black}{\Rightarrow 2 = x - 5 }\)

OpenStudy (anonymous):

Yes,thank you so much!!!!!! I a doing my brothers homework and have no clue how to do it!!!! But you guys are helping me understand!!!